State the null hypothesis and the alternative hypothesis in (a) words and (b) symbols for the following


A random sample of 200 students got a mean score of 62 with a standard deviation score of 5 in a knowledge test in mathematics. In the standardization of the test, μ= 50 with σ= 10


1
Expert's answer
2021-05-17T08:12:59-0400

a.)The null hypothesis is a typical statistical theory which suggests that no statistical relationship and significance exists in a set of given single observed variable, between two sets of observed data and measured phenomena.

b.)We have given that,

"n = 200"

"\bar x = 62"

"s = 5"

"\mu = 50"

"\sigma = 10"

Null hypothesis

"H_0 : \mu = 50"

Alternative hypothesis,

"H_1 : \mu \ne 50"

"z value = \dfrac{\bar x - \mu}{10} = \dfrac{62-50}{10} = 1.2"

At 0.05 level of significance z = 1.96

Hence, calculated value of z score is less than 1.96.

Null hypothesis is verified.